OpenStudy (anonymous):

rewrite as a logarithmic equation: 2^4=(a-b)

OpenStudy (anonymous):

take log for both side

OpenStudy (anonymous):

log2^4=log(a-b) 4log2=log(a-b)

OpenStudy (anonymous):

\[\log_{2} (a-b)=4\]

OpenStudy (anonymous):

Thanks so much. Would it be \[4=2\log(a-b)\]?

OpenStudy (anonymous):

it'd be \[\log_{2} \] almost certain

OpenStudy (anonymous):

Gottcha. Thank you so much.